What is the x-coordinate of the vertex for the graph of the equation y=2/3x^2-6x+4?
a) 1/18
b)1/3
c) 3
d) 9/2/2

Answer by stanbon(75887)

What is the x-coordinate of the vertex for the graph
of the equation y=2/3x^2-6x+4?
---
You have a quadratis with a = 2/3; b = -6; c = 4
----
The vertex occurs where x = -b/2a = 6/(4/3) = 18/4 = 9/2
================
Cheers,
Stan H.
================

a) 1/18
b)1/3
c) 3
d) 9/2/2